Badger Company is a job-order costing manufacturer that applies overhead on the basis of direct labor hours.

At the beginning of the year, to establish a predetermined overhead rate, Badger Company estimated total $4,500 in overhead costs and total 300 direct labor hours.

On March 1, there were two jobs in process: Job 89 with a beginning balance of $2,300 and Job 90 with a beginning balance of $1,824. During March, three new jobs were started: Jobs 91, 92, and 93. The amount of direct labor hours actually worked on each job during March was as follows:  

      Job 89   Job 90   Job 91   Job 92 Job 93

Direct labor hours    30 hrs.   50 hrs.   65 hrs.   25 hrs. 20 hrs.

Data on costs added during March for all jobs are as follows:  

      Job 89 Job 90   Job 91   Job 92   Job 93

Direct materials cost   $1,300 $1,726 $2,230    $1,525   $820

Direct labor cost       800    1,150    1,620        975      480

By March 31, Jobs 89 and 91 were completed and sold. Job 90 was completed but not sold yet. The rest of the jobs remained in process. On March 1, the beginning balance of the Finished Goods inventory account was $2,050. Badger Company prices its jobs at cost plus 40%.

1. What is the manufacturing overhead applied to Job 89?

A) $450

B) $750

C) $7,200

D) $12,000

2. What is the total manufacturing cost incurred by Job 90?

A) $3,626

B) $2,876

C) $5,450

D) $4,700

3. What is the balance in Work-In-Process on March 31?

A) $3,800

B) $2,875

C) $4,475

D) $9,925

4. What is the balance in Finished Goods on March 31?

A) $5,450

B) $7,500

C) $5,676

D) $17,175

5. What is the amount of Cost of Goods Sold for March?

A) $7,375

B) $9,675

C) $4,825

D) $10,300


6. What is the selling price of Job 91?

A) $6,755

B) $4,825

C) $1,930

D) $5,390
Answer #1

Overhead rate = Estimated overhead/estimated direct labor hours = 4500/300 = 15 per hour

1) Manufacturing overhead applied to Job 89 = 30*15 = $450

So answer is a) $450

2) Total manufacturing cost incurred by Job 90 = (1726+1150+50*15) = 3626

So answer is a) $3626

3) Balance in Work in process march 31 = (1525+820+975+480+45*15) = 4475

So answer is c) $4475

4) Balance in Finished goods march 31 = (3626+1824) = 5450

So answer is a) $5450

5) Amount of cost of goods sold = 1300+2230+800+1620+(95*15)+2300 = 9675

So answer is b) $9675

6) Selling price = 4825*1.4 = $6755

So answer is a) $6755